Calculate the magnitude of the faintest object that can be detected by an optical
telescope of diameter 5.0 m.
1
Expert's answer
2019-05-06T10:33:45-0400

For this problem we can use formula

"m=2+5lg(D)"


D - diameter telescope in millimeters; m - magnitude

Calculating

"m=2+5lg(5000)=20.5"